OpenStudy (kittykr98):

What is the equation of the parabola, in vertex form, with vertex at (2,-4) and directrix y=-6? x^2=12y answer choices: (y+4)^2=-4(x+2) (x-2)^2=-1/4(y+5) (x-2)^2=4(y+5) (y+2)^2=1/4(y+4)

OpenStudy (tkhunny):

What say you? (2,-4) implies something that looks like (x-2) an (y - (-4)) = (y+4) What can we throw out based on that observation?

OpenStudy (kittykr98):

you lost me with that second half

OpenStudy (kittykr98):

i think im supposed to convert this to standard form first? which is y=x^2 -12

OpenStudy (kittykr98):

i shouldve included the equation. x^2=12y

OpenStudy (kittykr98):

actually i have posted the wrong thing completely, sorry. i will tag you in the next thing

OpenStudy (tkhunny):

That equation appears to have nothing to do with the question. Here are the choices. A: (y+4)^2=-4(x+2) B: (x-2)^2=-1/4(y+5) C: (x-2)^2=4(y+5) D: (y+2)^2=1/4(y+4) <== This makes no sense. 'x' is missing. If the x-coordinate of the vertex is 2, we must have something that looks like "(x-2)". This eliminates A and D and we are left with two choices B: (x-2)^2=-1/4(y+5) C: (x-2)^2=4(y+5) If the y-coordinate of the vertex is -4, we must have something that looks like "(y+4)". This eliminates B and C and we are left with NO choices. Either you copied the problem incorrectly or the correct answer is "E: None of the Above". We didn't even use the Directrix information.